OpenStudy (madhu.mukherjee.946):

calculate the arithmetic mean: age group: 14-15 16-17 18-20 21-24 25-29 30-34 35-39 no.of wage earners: 60 140 150 110 110 100 90

OpenStudy (madhu.mukherjee.946):

its difficult to calculate the mid value here coz the common difference is changing

Nnesha (nnesha):

hmm don't we hve to add all the numbers and divide by the number of numbers

Nnesha (nnesha):

tag ganeshie8 or freckles

Nnesha (nnesha):

@madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

@ganeshie8 @freckles

OpenStudy (freckles):

@Nnesha is right

OpenStudy (freckles):

just add the numbers and divide by the number of numbers

OpenStudy (madhu.mukherjee.946):

no in group frequency we don't solve it like that way

OpenStudy (freckles):

oh I see the mid value is just the sum of endpoints and that result divided by 2

OpenStudy (freckles):

for example: \[\frac{14+15}{2}=\frac{29}{2}=14.5\]

OpenStudy (freckles):

so x_1=14.5 and f_1=60 x_1*f_1=14.5*60

OpenStudy (freckles):

\[\text{ mean }=\frac{x_1 f_1 +x_2 f_2 +x_3 f_3+x_4f_4+x_5f_5+x_6f_6+x_7f_7}{f_1+f_2+f_3+f_4+f_5+f_6+f_7} \\ \]

OpenStudy (freckles):

at least that is my definition of midpoint http://www.yale.edu/ynhti/curriculum/units/1986/5/86.05.03.x.html here they pick the midpoint of 99 to 100 as 98 which is odd because that is not the midpoint of 99 to 100

OpenStudy (madhu.mukherjee.946):

thank you so much

OpenStudy (freckles):

@madhu.mukherjee.946 still trying to figure out how to pick h  so I think I will choose a to be that thingy there because it looks like they are picking a to be one of the "middle numbers" and we only have one middle here since we have an odd amount of data entries do you ever seen this method? And do you know how to pick h?